Staggert Corp. will pay dividends of $5.00, $6.25, $4.75, and $3.00 in the next four years. Thereafter, the company expects its dividend growth rate to be constant at 9.0 percent. If the required rate of return is 19.5 percent, what is the current value of the stock?
Answer #1

Current value of the stock is $ 28.09

Working:

As per dividend discount model,
Current price of stock is the present value of future dividends.
Step-1:Present value of next four years dividend
Year dividend Discount factor Present value
a b c=1.195^-a d=b*c
1 $       5.00      0.8368 $       4.18
2 $       6.25      0.7003 $       4.38
3 $       4.75      0.5860 $       2.78
4 $       3.00      0.4904 $       1.47
Total $    12.82
Step-2:Terminal value of dividend
Terminal value of dividend = (D4*(1+g)/(Ke-g))*DF4 Where,
= (3*(1+0.09)/(0.1950-0.09))*0.4904 D4 $       3.00
= $    15.27 g 9%
Ke 19.50%
DF4      0.4904
Step-3:Present value of total dividends
Present value of total dividends = $    12.82 + $    15.27
= $    28.09
Thus,
Current value of stock is $    28.09